Convert each part of the ratio to percentage:    2: 3: 5 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

We have to find total parts by adding the given ratio = 2 + 3 + 5 = 10
1stpart = 2/1
= (2/10) × 100 %
= 2 × 10%
= 20%
2ndpart = 3/10
= (3/10) × 100%
= 3 × 10
= 30%
3rdpart = 5/10
= (5/10) × 100%
= 5 × 10
= 50%

Thanks